# Tomographic reconstruction of a disease transmission landscape via GPS recorded random paths

### v3

@@ 2.3 Testing
tests $H_{0}$ at level $\alpha$. One expects the TV-test to be more powerful than the likelihood ratio test for smooth (spatially homogeneous) alternative hypotheses (Sardy et al., 2022). A Monte Carlo simulation confirms it in Appendix B. As a side product, when the test is not rejected for some level $\alpha$ means that the solution to (2) for $\lambda=\lambda^{\rm QUT}_{\alpha}$ is a constant propensity map without having to solve (2).
@@ 2.4 Data augmentation with uncertainty quantification
GPS sampling constraints often result in small sample sizes $n$. We address it by performing data augmentation, however, introducing additional variability. Furthermore, the estimation of the propensity $\hat{{\boldsymbol{\mu}}}$ alone provides limited practical value for ecologists and practitioners. The estimate is conditional on the observed individuals and their random movements across the study area, which also introduces significant uncertainty. To account for the uncertainty arising from the randomness of the observed independent individuals and their limited sample size, we employ the bootstrap augmenting the sample size to $n_{\text{boot}}>n$ individuals. To generate spatially correlated tracks so as to reflect the gregarious behavior typical of herds, we sample the random moves for each resampled individual. This bootstrapped data augmentation procedure is repeated multiple times, and the pointwise $(1-\alpha)$ quantiles of the resulting bootstrapped propensity estimates provide a measure of uncertainty. However, the universal threshold $\lambda^{\rm QUT}_{\alpha}$ used to select the regularization parameter $\lambda$ tends to be conservative, shrinking the total variation differences of the propensity estimates toward zero and introducing bias. To improve both the estimation accuracy and the coverage properties of the uncertainty intervals, we apply a pointwise bias correction $\hat{\boldsymbol{\mu}}^{\text{bc}}=2\hat{\boldsymbol{\mu}}-\overline{\boldsymbol{\mu}}^{\text{boot}}$, where $\overline{\boldsymbol{\mu}}^{\text{boot}}$ represents the averaged bootstrapped propensity estimates (Efron & Tibshirani, 1993). We apply the data augmentation and uncertainty quantification in to the Cervid dataset using $n_{boot}=600$ and $1000$ sampled locations for each bootstrap individual with 100 runs.
@@ 3.1 Estimation of disease propensity maps
We generate an entire population of $n_{0}=5000$ individuals moving over a square fenced area $\Omega_{0}$ made of $N_{0}\times N_{0}$ lattice with known propensity ${\boldsymbol{\mu}}_{0}$ given by one of the three binary profiles in Figure 2 which simulates (from left to right) a lake, a river or a lake plus a corner. We choose a full discretization with $N_{0}=50$ to provide a fine mapping of the area. For individual $i\in\{1,\ldots,n_{0}\}$, we simulate $T=2880$ random moves $l^{(i)}_{1},l^{(i)}_{2},\cdots,l^{(i)}_{T}$ on the lattice $\Omega$. The value of $T=2880$ mimics GPS samplings every $t=15$ minutes for 30 days. At each time, the proposed moves are either to stay at the current location or to move to one of its four neighbors. We simulate two different herd behaviors: one half of the individuals moves randomly with equal probability $1/5$ to one of the five proposed moves, while the other half moves with a probability that depends on ${\boldsymbol{\mu}}_{0}$ in a way that doubles the probability of moving to a region of higher virus infection propensity. We call $L_{0}$ the $n_{0}\times T$ matrix which entry $(i,j)$ corresponds to the lattice locations of the $i$-th individual at time $t$. Based on $L_{0}$, we build the $n_{0}\times p_{0}$ matrix $X_{0}$ which counts the time spent in each of the $p_{0}=N_{0}^{2}$ lattice locations of $\Omega$ for each of the $n_{0}$ individuals. Then we generate the binary infection vector ${\bf y}_{0}$ whether the animal was infected during the time period, according to model (1). We expect more ones in the entries of ${\bf y}_{0}$ corresponding to the second herd which tends to stay in a region of higher virus infection propensity.
@@
For increasing sample sizes $n\in\{500,5000\}$, increasing spatial discretization $N\in\{30,50\}$, increasing temporal frequencies $1/t$ with $t\in\{96,1\}$, we perform Monte Carlo simulations for each combination of the three propensity maps ${\boldsymbol{\mu}}_{0}$ plotted in Figure 2. Each scenario is simulated 100 times, leading to 100 full data information triplets $({\bf y}_{0},L_{0},X_{0})$, from which a response vector ${\bf y}$ is extracted and a regression matrix $X$ is built, depending on the value of $(n,N,t)$. For each of the 100 data sets generated, we estimate the propensity map with our total variation estimator solution to (2) with the quantile universal threshold of Definition 2.2 for the value of $\lambda$. For comparison we also estimate the propensities using three approaches: splines using mgcv R package (Wood, 2001), a naive empirical estimate
$\hat{\mu}^{\rm emp}_{m}(\omega^{j,k})=\frac{\sum_{\{i|y_{i}=1\}}x_{i}^{j,k}}{\sum_{i=1}^{n}x_{i}^{j,k}}$ (5)
and its smoothed version using Gaussian process regression (GPR) on the log-odds
Regularization parameters for splines and GPR were chosen with cross-validation. Owing to the different values of $(n,N,t)$, we scale both the total variation, the splines, the empirical, and the GPR estimates on the interval $[0,1]$ to compare them in terms of mean squared error calculated as ${\rm MSE}=\sum_{l=1}^{p}(\hat{\mu}_{l}-\tilde{\mu}_{0,l})^{2}/p$, where $\tilde{\boldsymbol{\mu}}_{0}$ is the nearest-neighbor interpolation of ${\boldsymbol{\mu}}_{0}$ on a coarser scale and $p=N^{2}$.
Table 1 reports the MSE between $\hat{\boldsymbol{\mu}}$ and $\tilde{\boldsymbol{\mu}}_{0}$ averaged over the 100 Monte Carlo runs. Figure 3 shows typical estimations for the two propensity maps considered (river and lake+corner). Our method outperforms in nearly all scenarios. Splines tend to underestimate high propensity regions. As expected, MSE decreases for the asymptotic we considered, namely when the sample size $n$ of individuals increases, the spatial discretization $N$ increases, and the GPS frequency $1/t$ increases.
Figure 3: Increasing performance based on $(n,N,t)$ from $(500,30,96)$ to $(5000,50,1)$ for all estimators, river and lake+corner propensity maps.
Table 1: Results of Monte Carlo simulation reporting the averaged mean squared errors between $\hat{\boldsymbol{\mu}}$ (TV, splines, GPR and empirical) and ${\boldsymbol{\mu}}_{0}$ (${\tt lake}$, ${\tt river}$ and ${\tt lake+corner}$) over 100 runs for combinations of $(n,N,t)$. In bold the best estimator for every setup.
@@
N $t$ estimator lake river lake+corner
$30$ 96 TV 0.165 0.123 0.367 0.392 0.228 0.217
@@
GPR 0.421 0.360 0.541 0.471 0.439 0.385
empirical 0.709 0.460 0.740 0.545 0.729 0.491
1 TV 0.163 0.120 0.363 0.380 0.226 0.208
@@ 3.2 Coverage probability
We assess coverage probability of the bootstrap-based data augmentation and uncertainty quantification method outlined in Section 2.4 using the parameters $(n,N,t)=(500,30,1)$, with $n_{\text{boot}}=5000$ bootstrap individuals and 720 sampled locations for each bootstrap individual. A Monte Carlo simulation with 100 iterations, summarized in Table 2, evaluates the (MSE) of the bootstrap estimator, alongside the average coverage proportion and the average width of the point-wise confidence intervals. Targeting a nominal coverage level of 0.95, the total variation (TV) method exhibits reasonable coverage. Other methods show notably lower coverage. Compared with same $(n,N,t)$ setting in Table 1, the bootstrap estimate considerably improves the MSE for all estimators and simulated maps. Overall, the TV method demonstrates good performance with better coverage while maintaining reasonably sized confidence intervals. This suggests that our method provides a useful framework for practitioners to assess the significance of key propensity regions. To illustrate the uncertainty measures, Figure 4 presents typical point-wise uncertainty intervals for the four estimators based on the lake simulated profile.
Figure 4: Estimated profiles and point-wise uncertainty quantification lower and upper bounds of estimated lake propensity maps for all estimators, for fixed $(n,N,t)=(500,30,1)$, $n_{\text{boot}}=5000$ and 720 sampled locations for each bootstrapped individual.
Table 2: Results of Monte Carlo simulation reporting the averaged mean squared errors between $\hat{\boldsymbol{\mu}}^{\text{bc}}$ and ${\boldsymbol{\mu}}_{0}$, the averaged coverage probability (for a target value of $0.95$) and the averaged width of the point-wise bootstrap confidence intervals over 100 runs, for four estimators $\hat{\boldsymbol{\mu}}$ (TV, splines, GPR and empirical); for the three propensity maps ${\boldsymbol{\mu}}_{0}$ (${\tt lake}$, ${\tt river}$ and ${\tt lake+corner}$) for $(n,N,t)=(500,30,1)$, $n_{\text{boot}}=5000$, and 720 sampled locations for each bootstrap individual.
Average MSE, coverage and width
@@
estimator MSE coverage width MSE coverage width MSE coverage width
TV 0.036 0.843 0.322 0.215 0.778 0.820 0.072 0.817 0.500
splines 0.124 0.288 0.426 0.146 0.265 0.418 0.124 0.225 0.353
GPR 0.201 0.313 0.324 0.278 0.331 0.358 0.215 0.384 0.356
empirical 0.222 0.274 0.346 0.297 0.276 0.366 0.233 0.358 0.384
@@ 4 Application to white-tailed deer data
We now go back to the data described in Section 1.2 and we apply the discussed method. Figure 5 displays the estimated profiles obtained using our TV approach along with the associated point-wise confidence intervals (left panel). The right panel presents the observed abundance of C. stellifer, a known vector of EHDV. Figure 6 compares the EHDV-1 propensities estimated with four different methods.
Compared to the other models of disease risk estimation, TV points to smaller areas in which surveillance should be intensified for EHDV. It effectively pinpoints two specific feeders associated with EHDV-1 risk, despite having no prior knowledge of their locations. These two feeders have particularly high relative abundance of both Parous and Gravid C. stellifer which supports the risk estimation. In comparison, GPR and the empirical models suggest that the only area where disease risk is low is the northeastern region of the ranch. Such estimates make unlikely the surveillance or control efforts since that would require controlling the abundance of C. stellifer on 140 ha. Instead, the TV model suggests that the efforts should focus on the southeastern section and two supplementation feeders in the western portion. This reduces the initial control and surveillance area substantially. The splines model provided the most similar results to TV, albeit a larger risk area. Nonetheless, the splines model correlated well with the abundance of both Gravid and Parous C. stellifer.
@@
Figure 5: EHDV-1, EHDV-2 and EHD-6 disease risk estimation in white tailed deer and three exotic cervid species in a Florida ranch. The left column panels show the lower bound, estimate and upper bound of the relative disease risk for each of the three serotypes evaluate based on TV model. Warmer colors show higher risk of getting infected. The right panels show the average relative abundance of Culicoides stellifer in week 14 of 2015 and 2016. C. stellifer is a competent vector of EHDV.
Figure 6: EHDV-1 disease risk estimation in white tailed deer and three exotic cervid species in a Florida ranch, using four different methods.

### v4

@@ 3.3 Testing
tests $H_{0}$ at level $\alpha$. One expects the TV-test to be more powerful than the likelihood ratio test for smooth (spatially homogeneous) alternative hypotheses (Sardy et al., 2022). A Monte Carlo simulation confirms (See Supplementary materials). As a side product, when the test is not rejected for some level $\alpha$ means that the solution to (2) for $\lambda=\lambda^{\rm QUT}_{\alpha}$ is a constant propensity map without having to solve (2).
@@ 3.4 Data augmentation with uncertainty quantification
Estimation of the propensity alone provides limited practical value for ecologists and practitioners. Moreover GPS sampling constraints often result in small sample sizes $n$. To account for the uncertainty arising from the randomness of the observed independent individuals and their limited sample size, we employ the bootstrap that augments the sample size to $n_{\text{boot}}>n$ individuals. To generate spatially correlated tracks so as to reflect the gregarious behavior typical of herds, we sample the random moves for each resampled individual. This bootstrapped data augmentation procedure is repeated multiple times, and the pointwise $(1-\alpha)$ quantiles of the resulting bootstrapped propensity estimates provide a measure of uncertainty.
Also, due to the fact that the universal threshold $\lambda^{\rm QUT}_{\alpha}$ tends to be conservative, it shrinks and biases the total variation differences of the propensity estimates toward zero. To improve both the estimation accuracy and the coverage properties of the uncertainty intervals, we apply a pointwise bias correction $\hat{\boldsymbol{\mu}}^{\text{bc}}=2\hat{\boldsymbol{\mu}}-\overline{\boldsymbol{\mu}}^{\text{boot}}$, where $\overline{\boldsymbol{\mu}}^{\text{boot}}$ represents the averaged bootstrapped propensity estimates (Efron & Tibshirani, 1993). We perform uncertainty quantification on the Cervid dataset in Section 6.
4 Alternative methods
4.1 Direct preprocessing
An alternative data aggregation strategy computes a naive empirical estimator of the infection propensity at each spatial location:
$\hat{\mu}^{\rm emp}_{m}(\omega^{j,k})=\frac{\sum_{\{i|y_{i}=1\}}x_{i}^{j,k}}{\sum_{i=1}^{n}x_{i}^{j,k}}.$ (5)
This estimator can be interpreted as the empirical infection rate at each location, conditional on exposure. One may then smooth the log-odds transformation of this empirical rate:
Thanks to the structure induced by the aggregation, this approach enables the direct application of standard spatial smoothing methods.
4.2 Smoothing methods
Our data transformation casts the estimation problem as a tomographic linear inverse problem, that we regularized with total variation (Rudin et al., 1992). Whether we use the tomographic preprocessing data from Section 2 or the direct preprocessing data from Section 4.1, a range of other regularization strategies can be employed, depending on prior assumptions about spatial smoothness or structure. For example, splines- or wavelets-based smoothing (Wahba, 1990; Sardy et al., 2004) and Laplacian (Tikhonov) regularization (Zhou et al., 2003; Tikhonov, 1963) have a smoothness penalty. Spatially varying coefficients models (Gelfand et al., 2003) allow the effect of covariates or exposure to vary continuously across space, capturing local heterogeneity more flexibly. Other options include alternative penalties to accommodate transitions (Tibshirani et al., 2005), Gaussian Markov random fields (Rue & Held, 2005) for probabilistic spatial smoothing, or low-rank basis expansions for dimensionality reduction (Cressie & Johannesson, 2008), and Gaussian Process Regression (GPR) or kriging (Matheron, 1963).
@@ 5.1 Estimation of disease propensity maps
We generate an entire population of $n_{0}=5000$ individuals moving over a square fenced area $\Omega_{0}$ made of $N_{0}\times N_{0}$ lattice with known propensity ${\boldsymbol{\mu}}_{0}$ given by one of the three binary profiles in Figure 2 which simulates (from left to right) a lake, a river or a lake plus a corner. We choose a full discretization with $N_{0}=50$ to provide a fine mapping of the area. For individual $i\in\{1,\ldots,n_{0}\}$, we simulate $T=2880$ random moves $l^{(i)}_{1},l^{(i)}_{2},\cdots,l^{(i)}_{T}$ on the lattice $\Omega$. The value of $T=2880$ mimics GPS samplings every $t=15$ minutes for 30 days. At each time, the proposed moves are either to stay at the current location or to move to one of its four neighbors. We simulate two different herd behaviors: one half of the individuals moves randomly with equal probability $1/5$ to one of the five proposed moves, while the other half moves with a probability that depends on ${\boldsymbol{\mu}}_{0}$ in a way that doubles the probability of moving to a region of higher virus infection propensity. We call $L_{0}$ the $n_{0}\times T$ matrix which entry $(i,j)$ corresponds to the lattice locations of the $i$-th individual at time $t$. Based on $L_{0}$, we build the $n_{0}\times p_{0}$ matrix $X_{0}$ which counts the time spent in each of the $p_{0}=N_{0}^{2}$ lattice locations of $\Omega$ for each of the $n_{0}$ individuals. Then we generate the binary infection vector ${\bf y}_{0}$, whether or not the animals were infected during the time period, according to model (1). We expect more ones in the entries of ${\bf y}_{0}$ corresponding to the second herd which stays longer in regions of higher virus infection propensity.
@@
For increasing sample sizes $n\in\{500,5000\}$, increasing spatial discretization $N\in\{30,50\}$, increasing temporal frequencies $1/t$ with $t\in\{96,1\}$, we perform Monte Carlo simulations for each combination of the three propensity maps ${\boldsymbol{\mu}}_{0}$ plotted in Figure 2. Each scenario is simulated 1000 times, leading to 1000 full data information triplets $({\bf y}_{0},L_{0},X_{0})$, from which a response vector $\bf y$} is extracted and a regression matrix $X$ is built, depending on the value of $(n,N,t)$. For each of the 1000 data sets generated, we estimate the propensity map with our tomographic TV solution to (2) with the quantile universal threshold of Definition 3.2 for the value of $\lambda$. For comparison we also estimate the propensities using four smoothing approaches applied to our tomographic preprocessing data from Section 2 and four smoothing approaches applied to the alternative direct preprocessing data from Section 4.1. For the tomographic data we use Laplacian regularization with cross-validation, and splines and spatially variable coefficients (SVC) using mgcv R package (Wood, 2001). For the direct data we compare with the naive empirical estimate, and its smoothed versions using GPR with cross-validation, and total variation with tvR R package (Lee et al., 2022)) and Laplacian smoothing with regularization parameter $\lambda=\hat{\sigma}\log{p}$. Spatially varying coefficients and splines can be seen as equivalent to GPR for the direct preprocessing data.
Owing to the different values of $(n,N,t)$, we scale all estimates on the interval $[0,1]$ to compare them in terms of mean squared error calculated as ${\rm MSE}=\sum_{l=1}^{p}(\hat{\mu}_{l}-\tilde{\mu}_{0,l})^{2}/p$, where $\tilde{\boldsymbol{\mu}}_{0}$ is the nearest-neighbor interpolation of ${\boldsymbol{\mu}}_{0}$ on a coarser scale and $p=N^{2}$.
Table 1 reports the MSE between $\hat{\boldsymbol{\mu}}$ and $\tilde{\boldsymbol{\mu}}_{0}$ averaged over the 1000 Monte Carlo runs. Figure 3 shows typical estimations for the river propensity map. Our tomographic TV method outperforms in nearly all scenarios. Among the tomographic estimators, spline‑based and SVC reconstructions consistently underestimate the magnitude of high‑propensity regions, whereas the Laplacian method fails to detect these areas altogether. Although the direct smoothing estimators appear slightly crisper with larger sample sizes, they still yield the highest mean‑squared error and produce increasingly noisy propensities. As expected, MSE decreases for the asymptotic we considered, namely when the sample size $n$ of individuals increases, the spatial discretization $N$ increases, and the GPS frequency $1/t$ increases.
Figure 3: Performance based on increasing $(n,N,t)$ from $(500,30,96)$ to $(5000,50,1)$ for all tomographic smoothing estimators (left) and direct smoothing estimators (right), for river propensity map.
Table 1: Results of Monte Carlo simulation reporting the averaged mean squared errors between $\hat{\boldsymbol{\mu}}$ for four tomographic smoothing methods and four direct smoothing methods, and ${\boldsymbol{\mu}}_{0}$ ($\tt lake$}, $\tt river$} and $\tt lake+corner$}) over 1000 runs for combinations of $(n,N,t)$. In bold the best estimator for every setup.
@@
N $t$ data estimator lake river lake+corner
$30$ 96 tomographic TV 0.165 0.123 0.367 0.392 0.228 0.217
@@
Laplacian 0.351 0.234 0.437 0.307 0.356 0.249
SVC 0.225 0.193 0.345 0.309 0.240 0.222
direct TV 0.232 0.205 0.525 0.334 0.331 0.229
GPR 0.421 0.36 0.541 0.471 0.439 0.385
Laplacian 0.549 0.411 0.674 0.578 0.629 0.526
empirical 0.709 0.461 0.74 0.545 0.729 0.491
30 1 tomographic TV 0.163 0.120 0.363 0.380 0.226 0.208
@@ 5.2 Coverage probability
We assess coverage probability of the bootstrap-based data augmentation and uncertainty quantification method outlined in Section 3.4 using the parameters $(n,N,t)=(500,30,1)$, with $n_{\text{boot}}=5000$ bootstrap individuals and 720 sampled locations for each bootstrap individual. A Monte Carlo simulation with 1000 iterations, summarized in Table 2, evaluates the average total coverage proportion of the point-wise confidence intervals. Since practitioners are mostly concerned with distinguishing high‑ and low‑risk areas, we also report the average coverage proportion separately for the low‑risk and high‑risk regions. Targeting a nominal coverage level of 0.95, the tomographic TV method exhibits close to nominal coverage for the high-risk region, while reasonable coverage for low-risk region. Other methods show notably lower coverage, with methods applied with the direct preprocessing being the worst. Overall, the tomographic TV method demonstrates good performance with better coverage. The diminished coverage in low‑risk regions is unsurprising, reflecting the shrinkage bias typical of penalized estimators. In contrast, the near‑nominal coverage achieved in high‑risk regions confirms that the areas of greatest concern are accurately detected. This suggests that our method provides a useful framework for practitioners to assess the significance of critical propensity hotspots. To illustrate the uncertainty measures, Figure 4 presents typical point-wise uncertainty intervals for the eight estimators based on the lake simulated profile.
Figure 4: Estimated profiles and point-wise uncertainty quantification lower and upper bounds of estimated lake propensity map for tomographic smoothing estimators (left) and direct smoothing estimators (right), for fixed $(n,N,t)=(500,30,1)$, $n_{\text{boot}}=5000$ and 720 sampled locations for each bootstrapped individual.
Table 2: Results of Monte Carlo simulation reporting the averaged low-risk coverage, averaged high-risk coverage and averaged overall coverage probability (for a target value of $0.95$) of the point-wise bootstrap confidence intervals over 1000 runs for $\hat{\boldsymbol{\mu}}$ for four tomographic smoothing estimators and four direct smoothing estimators; for the three propensity maps ${\boldsymbol{\mu}}_{0}$ ($\tt lake$}, $\tt river$} and $\tt lake+corner$}), $(n,N,t)=(500,30,1)$, $n_{\text{boot}}=5000$, and 720 sampled locations for each bootstrap individual. In bold the closest to the nominal level.
Average low-risk, high-risk and total coverage
@@
estimator low high global low high overall low high overall
tomographic TV 0.83 0.95 0.85 0.73 0.99 0.79 0.83 0.93 0.84
splines 0.26 0.60 0.30 0.22 0.44 0.26 0.20 0.40 0.24
Laplacian 0.92 0.86 0.91 0.72 0.69 0.72 0.82 0.71 0.81
SVC 0.30 0.20 0.29 0.29 0.62 0.36 0.38 0.21 0.35
direct TV 0.66 0.84 0.69 0.55 0.97 0.63 0.71 0.87 0.73
GPR 0.33 0.45 0.34 0.29 0.45 0.32 0.39 0.43 0.40
Laplacian 0.23 0.37 0.24 0.08 0.42 0.15 0.09 0.32 0.13
empirical 0.29 0.34 0.30 0.26 0.34 0.28 0.37 0.32 0.37
@@ 6 Application to white-tailed deer data
Figure 5: EHDV-1, EHDV-2 and EHD-6 disease risk estimation in white tailed deer and three exotic cervid species in a Florida ranch. The left column panels show the lower bound, estimate and upper bound of the relative disease risk for each of the three serotypes evaluate based on tomographic TV estimator. Warmer colors show higher risk of getting infected. The right panels show the average relative abundance of Culicoides stellifer in week 14 of 2015 and 2016. C. stellifer is a competent vector of EHDV. Each map shows feeder locations (black dots) and water sources (bright light-blue shapes and lines).
We return to the data of Section 1.2 and apply the proposed method, incorporating data augmentation and uncertainty quantification using $n_{boot}=600$ and $1000$ sampled locations for each bootstrap individual with 1000 runs.
Figure 5 displays the estimated profiles obtained using our tomographic TV approach along with the associated point-wise confidence intervals (left panel). The right panel presents the observed abundance of C. stellifer, a known vector of EHDV. Figure 6 compares the EHDV-1 propensities estimated with four tomographic smoothing methods and four direct smoothing methods. Figures 5 and 6 display feeder locations (black dots) and water sources (bright light-blue shapes and lines) for context only; these features are not included as covariates in any of the models.
Compared to the other models of disease risk estimation, tomographic TV points to smaller areas in which surveillance should be intensified for EHDV. It effectively pinpoints two specific feeders associated with EHDV-1 risk, despite having no prior knowledge of their locations. These two feeders have particularly high relative abundance of both Parous and Gravid C. stellifer which supports the risk estimation. In comparison, most of the other models suggest that the only area where disease risk is low is the northeastern region of the ranch. Such estimates make unlikely the surveillance or control efforts since that would require controlling the abundance of C. stellifer on 140 ha. Instead, the tomographic TV model suggests that the efforts should focus on the southeastern section and two supplementation feeders in the western portion. This reduces the initial control and surveillance area substantially. The tomographic splines and SVC models provided the most similar results to tomographic TV, albeit a larger risk area. Nonetheless, the both models correlated well with the abundance of both Gravid and Parous C. stellifer.
@@
Figure 6: EHDV-1 disease risk estimation in white tailed deer and three exotic cervid species in a Florida ranch, using four tomographic smoothing estimators (first row) and four direct smoothing estimators (second row). Each map shows feeder locations (black dots) and water sources (bright light-blue shapes and lines).